# Physics-informed machine learning for combustion: A review

### v1

@@ 1 Introduction
As shown in Figure 1(a), combustion research and applications involve three core elements: physics, model, and data. Physics represents the real world; models aim to represent physics either through theoretical/physical formulations or data-driven approaches; and data are recorded values that typically contain physics. Transformations among these elements constitute the forward and inverse problems in research. Forward problems include physics-to-data and model-to-data problems, which involve realizing combustion processes in the real and virtual world, respectively, to generate data. The former includes experimental research, such as designing methods for controllable combustion 9, while the latter primarily refers to numerical simulations, including application to various combustion devices 10, 11, 12, 13 and the development of numerical algorithms 14, 15. In contrast, inverse problems include physics-to-model and data-to-model problems, which develop models or theories based on first principles 4, 5, 6, 7, 8 and data patterns, respectively. Model development can involve constructing models from scratch or determining parameters for existing models, which may take either interpretable or purely fitted forms. Forward problems and inverse problems are inherently interdependent. For example, the model-to-data forward problems require the complete form and parameters of the model to be known in advance, while data-to-model inverse problems often depend on physics-to-data experiments 16, 17, 18, 19. Consequently, theoretical, experimental, and numerical methods are often adopted simultaneously 20. However, these three types of methods differ substantially in implementation, resource requirements, cost, output form, and transferability, hindering their integration and slowing the research cycle. This highlights the need for a paradigm that can overcome these barriers, establish a unified framework linking physics, models, and data, and more effectively address a wide range of forward and inverse problems in combustion.
@@
Figure 1: (a) Physics, model, and data are three elements in combustion. Physics represents the real world, physical models and ML models try to represent physics, while data are value records and usually contain physics. Their transformations constitute the forward and inverse problems, which are defined as generating data and modeling, respectively. (b) PIML can act as a unified bridge connecting physics, model, and data. PIML is constituted by physical model, ML model, and possible data, then each of them can be enhanced by the other two. The dashed colored arrows represent optional implementations.
@@
Artificial intelligence (AI), machine learning (ML) 21, and deep learning (DL) 22 have made remarkable achievements in recent years, transforming fields such as computer vision (CV) 23, natural language processing (NLP) 24, gaming 25, and weather forecasting 26, 27, 28. AI for science (AI4Sci) and scientific machine learning (SciML) 29 have emerged as key areas of interest for both the AI and scientific communities. These approaches leverage AI/ML to analyze scientific data, uncover patterns 30, and solve scientific problems, and are recognized as the fourth paradigm of science 31, 32, 33. They have produced transformative results, such as the AlphaFold in life sciences 34, 35. SciML has also made significant advances in fluid mechanics 36, 37, 38, including both experimental fluid mechanics 39 and computational fluid dynamics (CFD) 40, 41, 42. The combustion community has likewise embraced ML 43, 44, motivated not only by the power and versatility of ML methods, but also by the data-rich nature of combustion and the limitations of traditional approaches. ML has been applied to combustion model optimization 45, 46, 47, 48, 49, combustion chemistry 50, 51, 52, 53, 54, combustion tabulation 55, 56, 57, 58, 59, and surrogate modeling of combustion systems 60, 61, 62, 63, 64. The applied ML techniques include supervised, unsupervised, and semi-supervised learning, which respectively focus on learning mappings from labeled data, identifying patterns in unlabeled data, and combining both approaches. These methods are built on diverse ML models, including neural networks (NNs), support vector machines (SVMs), and random forests (RFs). For more applications of ML for combustion, please refer to these reviews 43, 44. Data-driven ML has greatly advanced the data-to-model paradigm, enabling the exploitation of big data in combustion and the resolution of many longstanding challenges in the field.
@@ 2.2 Reviews and scoping
We first introduce the reviews on PIML in a chronological sequence. Karniadakis et al. 65 first provided a comprehensive review of PIML. They pointed out three ways to embed physics into machine learning: observational biases, inductive biases, and learning biases, which correspond to preprocessing the data, designing the model architecture, and modifying the training algorithm, respectively. The classical PINNs belong to the third category, which impose the physical constraints in the loss function, which is a soft manner. Inductive biases is another popular way, with CNNs being the most celebrated example, which respect the invariance in images. Other examples include encoding symmetries, PDEs 73, 74, and chemical laws 75 in model architectures. Shukla et al. 76 reviewed not only the classical PINNs but also how physics-informed learning can be accomplished with graph NNs (GNNs). Cuomo et al. 77 provided a comprehensive review of PINNs and their variants, covering the NN architectures, learning approaches, theories, applications, and software. Hao et al. 78 presented a survey on PIML from the perspective of machine learning researchers. The physics priors were categorized into three types: differential equations, symmetry constraints, and intuitive physics. Various approaches to incorporating physical prior into ML were identified: data, model, objective, optimizer, and inference. They categorized the PIML tasks into two classes: using PIML to solve scientific problems and to solve ML problems. For solving scientific problems, the neural simulation and inverse problems were introduced, where the neural simulation were classified into the neural solvers (i.e., PINNs) and neural operators (NOs). For solving ML problems, the PIML for computer vision and reinforcement learning was summarized. Wang et al. 79 provided an expert’s guide to training PINNs, presenting a series of best practices to improve the efficiency and accuracy of PINNs, such as non-dimensionalization, NN architectures, respecting temporal causality, loss balancing, and hard constraints. Faroughi et al. 80 presented a review of four NN frameworks: physics-guided NNs, physics-informed NNs, physics-encoded NNs 73, and NOs. The former three frameworks correspond to the observational, learning, and inductive biases in Reference 65, respectively. Kim et al. 81 reviewed the PINNs on both forward and inverse problems and gave special attention to the multiscale analysis using PINNs. Farea et al. 82 reviewed the techniques, applications, trends, and challenges of PINNs. The three ways of integrating prior knowledge into PINNs were pointed out: feature engineering, model construction, and regularization, which is consistent with the taxonomy in References 65, 80. Zhang et al. 83 summarized the limitations of PINNs and identified three root causes for their failure: poor multiscale approximation ability and ill-conditioning, weak mathematical rigor, and inadequate integration of physical information. Corresponding future directions and prospects were also provided. More recently, Toscano et al. 84 reviewed the recent advances in PIML. They identified three key components of the PIML framework: representation model, governing equation, and optimization process, which correspond to the blue, yellow, and green part of Figure 4, respectively, and is consistent with the taxonomy in Reference 85. They also reviewed the applications, uncertainty quantification (UQ), theoretical advances, and computational frameworks of PIML, and provided a chronological overview of key advancements in PIML. Meng et al. 86 surveyed a wide variety of recent works in PIML. They presented four categories of physics knowledge in PIML: (1) classical mechanics and energy conservation laws, (2) symmetry, invariant and equivariant functions, (3) numerical methods for PDEs, and (4) Koopman theory. Consistent with References 65, 80, 82, they also identified three ways to integrate physics knowledge into ML: data enhancement, NN architecture design, and optimization.
@@
Some reviews focus on the PIML for fluid mechanics 87, 88, 89, which can provide some insights for PIML for combustion flows. Note that the classification of PIML in Reference 88 is again consistent with the References 65, 80, 82, 86, including the physics-informed features and labels, architectures, and loss functions. However, these reviews rarely include combustion or reacting flows. Recently, Sophiya et al. 90 reviewed the PINNs in industrial gas turbines (IGTs) research, highlighting their diverse applications in aerodynamics and aeromechanics. However, there are few combustion applications introduced. A more combustion-related review is Reference 91, which focuses on the PIML in multiphysics modeling in chemical engineering. The applications of PIML in momentum transfer/fluid flow, heat transfer, mass transfer, and chemical reaction were first reviewed separately, then the PIML for multiphysics coupling, surrogate model construction, and design optimization were presented. PIML for chemical engineering has also been briefly reviewed in Reference 84. However, it should be noted that combustion and chemical engineering do not contain each other, but intersect with each other. They both involve chemical reactions but the range of combustion is more narrower: a high-temperature exothermic redox chemical reaction by definition. On the other hand, combustion is not limited to chemical engineering, but also involves power generation, transportation, fire safety, and others.
@@
In a broader context, the classical PIML belongs to the category of AI for PDEs (AI4PDEs), which also includes the operator learning. Some researchers have reviewed the AI4PDEs, including DL for PINN and related parameter identification problems 92, promising directions of AI4PDEs 93, and AI4PDEs in computational mechanics 94. Function learning approximates the mapping between finite-dimensional tensor spaces, while operator learning approximates the mapping between infinite-dimensional function spaces. A representative operator is the solution operator of PDEs, which maps the IC/BCs or parameters (e.g., operating, geometrical, physical) to the solution functions. Function learning can only learn a single PDE solution per training, so the model needs to be retrained once conditions or parameters change. As a progress, operator learning can learn parametric PDE solutions at once, providing a method for rapid PDE-based predictions in practice. When the model of operator learning is an NN, it is referred to as neural operator (NO). Two common NOs are the deep operator network (DeepONet) 95 and the Fourier neural operator (FNO) 96. However, they are both purely data-driven, so their physics-informed versions have also been proposed, i.e., the PI-DeepONet 97 and PI-FNO 98.
@@
Based on these reviews, some conclusions can be drawn. (1) The scope of PIML has been expanded from the classical PINNs to a broader sense. Table 1 compares the PIML in a narrow and in a broad sense from four aspects: model architecture, ways to inform physics, form of physics, and differentiation. In this paper, both types of PIML will be introduced. Unless otherwise specified, PIML in this paper refers to PIML in a broad sense, while PINNs refer to PIML using NNs as the representation model. Besides, some related terms, such as "physics-guided", "physics-embedded", "physics-encoded", "physics-enforced", "physics-constrained", and "physics-preserved", will all be viewed as "physics-informed" in a broad sense in this paper. (2) Many reviews have consensus on three ways to inform physics 65, 80, 82, 86, 88: through data, (ML) model, and optimization, as summarized in Table 2. Although the data is obtained by scientific means, ML based solely on physics-informed data is still purely data-driven, so it is a weak mechanism for informing physics, as stated in Reference 65. Therefore, this paper mainly introduces the latter two methods, and the first way will be introduced only if it has a strong embedding. (3) There is a lack of systematic reviews of PIML for combustion, which is the aim of this paper. (4) Many reviews introduce PIML and NOs together. However, most NOs are purely data-driven. This paper will only introduce PIML and physics-informed NOs (PINOs) in detail. By our definition, the latter can also be regarded as a kind of PIML.
@@
Model architecture MLP All models: CNN 99, 100, Transformer 101, DeepONet 97, 102, 103, 104, 105, 106, 107, 108, KAN 109, 110
How to inform physics Loss function All ways: Model architecture 73, 74, 75, input transformation, output correction 111, 112, 113, 51, 52, numerical solver 114, 115, 113, 116
Form of physics PDEs All forms: algebraic equation, inequality, known coefficients 117, 118, 119, 120
Differentiation AD All methods, such as ND: convolution kernels for space 99, 100, classical schemes for time 121, 73
@@ 2.3 Developments
This subsection will briefly introduce the developments of PIML, focusing mainly on PINNs that use ODE/PDEs as physical constraints, as this is the main and classical part of PIML. Most introductions are brief; for detailed introduction, please refer to the reviews in Section 2.2. Consistent with the taxonomy of References 84, 85, this paper divides PIML into three components: representation model, optimization process, and PDE system, which will be introduced separately in this subsection. Then some miscellaneous important topics will be introduced.
@@ 2.3.1 Representation model
Activation function. Activation functions ($\sigma$ in Figure 4) introduce nonlinearity to an NN. In order to implement AD to solve PDEs, the activation functions of PINNs need to have good differentiability. The $\tanh()$ and $\sin()$ are two common choices. Some adaptive activation functions have been proposed to accelerate the PINNs 134, 135.
@@
Architecture. Based on the representation way of the NN, PINNs can be divided into two categories: continuous and discrete PINNs, which are compared in Table 3. Continuous PINNs follow the coordinate-to-variable (query-to-value) mapping way, i.e., $\widehat{\mathbf{u}}=\mathrm{NN}(\mathbf{z})$, while discrete PINNs follow the field-to-field mapping way, i.e., $\widehat{\mathbf{U}}_{\mathrm{out}}=\mathrm{NN}(\mathbf{U}_{\mathrm{in}})$, where $\mathbf{U}_{\mathrm{in}}$ and $\widehat{\mathbf{U}}_{\mathrm{out}}$ are the input and output fields of the NN, respectively. Continuous PINNs provide continuous representations because they can directly predict the field values at any given query coordinates without interpolation, while discrete PINNs provide discrete representations because $\mathbf{U}_{\mathrm{in}}$ and $\widehat{\mathbf{U}}_{\mathrm{out}}$ are always discrete fields on meshes. When $\mathbf{U}_{\mathrm{in}}$ and $\widehat{\mathbf{U}}_{\mathrm{out}}$ are the fields defined on coarse and fine grids, respectively, the task is called super-resolution (SR) and has been applied to combustion (see Section 5.5). When $\mathbf{U}_{\mathrm{in}}$ and $\widehat{\mathbf{U}}_{\mathrm{out}}$ are fields at time $t$ and $t+\Delta t$, respectively, the task is usually referred to as autoregressive (AR) learning 136, 137. In this case, the NN can be called a time stepper, next-step predictor, forward propagator 138, or emulator 138, 137. On the contrary, the classical continuous PINNs belong to the paradigm of non-autoregressive (NAR) learning 136. The AR paradigm has been widely used in surrogate modeling of chemical kinetics. See Section 4.2 for details. Note that the AR and NAR learning can also be combined to form a hybrid paradigm 136, 139, 140. Instead of using AD in continuous PINNs, discrete PINNs employ numerical differentiation (ND) to compute derivatives. In other words, the PDE residual is $\mathcal{F}[\widehat{\mathbf{U}}_{\mathrm{out}}(\mathbf{U}_{\mathrm{in}})]$ computed on $\mathbf{U}$ using ND, instead of $\mathcal{F}[\widehat{\mathbf{u}}(\mathbf{z})]$ in Equation (2.5) computed on $\widehat{\mathbf{u}}$ and $\mathbf{z}$ using AD. For spatial derivatives, ND can be achieved using differential convolution kernels (similar to the gradient operators in computer vision 141), so convolutional NNs (CNNs) are usually adopted. For temporal derivatives, ND can be achieved via classical numerical schemes, such as the Euler scheme. In this case, recurrent NNs (RNN) can be used for modeling the temporal evolution. The choice between continuous and discrete PINNs will be discussed in Section 8.2.
@@
  • Continuous PINNs. The architectures of continuous PINNs mainly include the MLP, ELM 142, radial basis network (RBN) 143, PointNet 144, DeepONet 95, and Kolmogorov-Arnold network (KAN) 145, 146. Their physics-informed versions have been proposed: PI-ELM 147, PI-RBN 148, PI-PointNet 149, PI-DeepONet 97, and PI-KAN 109, 110. Moreover, the basic MLP has been modified in PINNs, including the multiple MLPs 150, modified MLP 79, and the physics-informed residual adaptive networks (PirateNets) 151, which tackles the issue that PINNs have difficulty using very deep NNs. Some generative models, such as the generative adversarial network (GAN) 152 and variational autoencoder (VAE) 153, can also use MLPs as the building blocks and be extended to the PI-GAN 154 and PI-VAE 155.
@@
  • Discrete PINNs. Basic architectures of discrete PINNs include the RNN (represented by the long short-term memory (LSTM) network 156), CNN 157, 23, recurrent CNN (RCNN), GNN 158, and Transformer 159. Some generative models usually use them as the backbone, such as the VAE, GAN, and diffusion model (DM) 160, 161, 162. The basic architectures can also serve as the building blocks of NOs, such as the Fourier neural operator (FNO) 96. Most of these NNs have their physics-informed extensions, including the PI-LSTM 163, PI-CNN 99, 100, PI-RCNN (or PhyCRNet) 121, PI-GNN 164, PI-Transformer 165, 101, PI-DM 166, and PI-FNO 98. Note that the architectures mentioned in continuous PINNs can also be used in discrete PINNs as long as the mapping is a field-to-field way. The same goes for the generative models, which are architecture-independent. See examples in Section 4.2 and Section 5.5. In fact, the discrete time models proposed in the original paper of PINN can be viewed as a discrete PINN on the temporal dimension 66, which adopted the MLP and Runge-Kutta scheme. However, the architectures just mentioned (such as the CNNs) are rarely applied to continuous PINNs because they typically do not take coordinates as network inputs.

### v2

@@ 1 Introduction
As shown in Figure 1(a), combustion research and applications involve three core elements: physics, model, and data. Physics represents the real world; models aim to represent physics either through theoretical/physical formulations or data-driven approaches; and data are recorded values that typically contain physics. Transformations among these elements constitute the forward and inverse problems in research. Forward problems include physics-to-data and model-to-data problems, which involve realizing combustion processes in the real and virtual world, respectively, to generate data. The former includes experimental research, such as designing methods for controllable combustion 9, while the latter primarily refers to numerical simulations, including application to various combustion devices 10, 11, 12, 13 and the development of numerical algorithms 14, 15. In contrast, inverse problems include physics-to-model and data-to-model problems, which develop models or theories based on first principles 4, 5, 6, 7, 8 and data patterns, respectively. Model development can involve constructing models from scratch or determining parameters for existing models, which may take either interpretable or purely fitted forms. Forward problems and inverse problems are inherently interdependent. For example, the model-to-data forward problems require the complete form and parameters of the model to be known in advance, while data-to-model inverse problems often depend on physics-to-data experiments 16, 17, 18, 19, 20. Consequently, theoretical, experimental, and numerical methods are often adopted simultaneously 21. However, these three types of methods differ substantially in implementation, resource requirements, cost, output form, and transferability, hindering their integration and slowing the research cycle. This highlights the need for a paradigm that can overcome these barriers, establish a unified framework linking physics, models, and data, and more effectively address a wide range of forward and inverse problems in combustion.
@@
Figure 1: (a) Physics, model, and data are three elements in combustion. Physics represents the real world, physical models and ML models try to represent physics, while data are value records and usually contain physics. Their mutual transformations constitute the forward and inverse problems, which are defined as generating data and modeling, respectively. These transformations also correspond to four basic paradigms of science, as labeled in the figure. (b) PIML can act as a unified bridge connecting physics, model, and data. PIML is constituted by physical model, ML model, and possible data, then each of them can be enhanced by the other two. Models can better represent physics, while data can better reflect physics. The dashed colored arrows represent optional implementations.
@@
Artificial intelligence (AI), machine learning (ML) 22, and deep learning (DL) 23 have made remarkable achievements in recent years, transforming fields such as computer vision (CV) 24, natural language processing (NLP) 25, gaming 26, and weather forecasting 27, 28, 29. AI for science (AI4Sci) and scientific machine learning (SciML) 30 have emerged as key areas of interest for both the AI and scientific communities. These approaches leverage AI/ML to analyze scientific data, uncover patterns 31, and solve scientific problems, and are recognized as the fourth paradigm of science 32, 33, 34. They have produced transformative results, such as the AlphaFold in life sciences 35, 36. SciML has also made significant advances in fluid mechanics 37, 38, 39, including both experimental fluid mechanics 40 and computational fluid dynamics (CFD) 41, 42, 43. The combustion community has likewise embraced ML 44, 45, 46, 47, motivated not only by the power and versatility of ML methods, but also by the data-rich nature of combustion and the limitations of traditional approaches. ML has been applied to combustion model optimization 48, 49, 50, 51, 52, combustion chemistry 53, 54, 55, 56, 57, combustion tabulation 58, 59, 60, 61, 62, and surrogate modeling of combustion systems 63, 64, 65, 66, 67. The applied ML techniques include supervised, unsupervised, and semi-supervised learning, which respectively focus on learning mappings from labeled data, identifying patterns in unlabeled data, and combining both approaches. These methods are built on diverse ML models, including neural networks (NNs), support vector machines (SVMs), and random forests (RFs). For more applications of ML for combustion, please refer to these reviews 44, 45, 46. Data-driven ML has greatly advanced the data-to-model paradigm, enabling the exploitation of big data in combustion and the resolution of many longstanding challenges in the field.
@@ 2.2 Reviews and scoping
We first introduce the reviews on PIML in a chronological sequence. Karniadakis et al. 68 first provided a comprehensive review of PIML. They pointed out three ways to embed physics into machine learning: observational biases, inductive biases, and learning biases, which correspond to preprocessing the data, designing the model architecture, and modifying the training algorithm, respectively. The classical PINNs belong to the third category, which impose the physical constraints in the loss function, which is a soft manner. Inductive biases is another popular way, with CNNs being the most celebrated example, which respect the invariance in images. Other examples include encoding symmetries, PDEs 76, 77, and chemical laws 78 in model architectures. Shukla et al. 79 reviewed not only the classical PINNs but also how physics-informed learning can be accomplished with graph NNs (GNNs). Cuomo et al. 80 provided a comprehensive review of PINNs and their variants, covering the NN architectures, learning approaches, theories, applications, and software. Hao et al. 81 presented a survey on PIML from the perspective of machine learning researchers. The physics priors were categorized into three types: differential equations, symmetry constraints, and intuitive physics. Various approaches to incorporating physical prior into ML were identified: data, model, objective, optimizer, and inference. They categorized the PIML tasks into two classes: using PIML to solve scientific problems and to solve ML problems. For solving scientific problems, the neural simulation and inverse problems were introduced, where the neural simulation were classified into the neural solvers (i.e., PINNs) and neural operators (NOs). For solving ML problems, the PIML for computer vision and reinforcement learning was summarized. Wang et al. 82 provided an expert’s guide to training PINNs, presenting a series of best practices to improve the efficiency and accuracy of PINNs, such as non-dimensionalization, NN architectures, respecting temporal causality, loss balancing, and hard constraints. Faroughi et al. 83 presented a review of four NN frameworks: physics-guided NNs, physics-informed NNs, physics-encoded NNs 76, and NOs. The former three frameworks correspond to the observational, learning, and inductive biases in Reference 68, respectively. Kim et al. 84 reviewed the PINNs on both forward and inverse problems and gave special attention to the multiscale analysis using PINNs. Farea et al. 85 reviewed the techniques, applications, trends, and challenges of PINNs. The three ways of integrating prior knowledge into PINNs were pointed out: feature engineering, model construction, and regularization, which is consistent with the taxonomy in References 68, 83. Zhang et al. 86 summarized the limitations of PINNs and identified three root causes for their failure: poor multiscale approximation ability and ill-conditioning, weak mathematical rigor, and inadequate integration of physical information. Corresponding future directions and prospects were also provided. More recently, Toscano et al. 87 reviewed the recent advances in PIML. They identified three key components of the PIML framework: representation model, governing equation, and optimization process, which correspond to the blue, yellow, and green part of Figure 4, respectively, and is consistent with the taxonomy in Reference 88. They also reviewed the applications, uncertainty quantification (UQ), theoretical advances, and computational frameworks of PIML, and provided a chronological overview of key advancements in PIML. Meng et al. 89 surveyed a wide variety of recent works in PIML. They presented four categories of physics knowledge in PIML: (1) classical mechanics and energy conservation laws, (2) symmetry, invariant and equivariant functions, (3) numerical methods for PDEs, and (4) Koopman theory. Consistent with References 68, 83, 85, they also identified three ways to integrate physics knowledge into ML: data enhancement, NN architecture design, and optimization.
@@
Some reviews focus on the PIML for fluid mechanics 90, 91, 92, which can provide some insights for PIML for combustion flows. Note that the classification of PIML in Reference 91 is again consistent with the References 68, 83, 85, 89, including the physics-informed features and labels, architectures, and loss functions. However, these reviews rarely include combustion or reacting flows. Recently, Sophiya et al. 93 reviewed the PINNs in industrial gas turbines (IGTs) research, highlighting their diverse applications in aerodynamics and aeromechanics. However, there are few combustion applications introduced. A more combustion-related review is Reference 94, which focuses on the PIML in multiphysics modeling in chemical engineering. The applications of PIML in momentum transfer/fluid flow, heat transfer, mass transfer, and chemical reaction were first reviewed separately, then the PIML for multiphysics coupling, surrogate model construction, and design optimization were presented. PIML for chemical engineering has also been briefly reviewed in Reference 87. However, it should be noted that combustion and chemical engineering do not contain each other, but intersect with each other. They both involve chemical reactions but the range of combustion is narrower: a high-temperature exothermic redox chemical reaction by definition. On the other hand, combustion is not limited to chemical engineering, but also involves power generation, transportation, fire safety, and others. Deng et al. 47 reviewed the SciML in combustion, presenting how SciML can assist: (1) modeling, such as in parameter estimation and reaction mechanism generation, (2) simulation, such as via DL and knowledge-guided ML, and (3) reconstruction, such as for sparse reconstruction and mutual inference. However, this review does not focus on PIML and does not elaborate on how physics and ML can be integrated in diverse ways, which are the main points of this paper.
@@
In a broader context, the classical PIML belongs to the category of AI for PDEs (AI4PDEs), which also includes the operator learning. Some researchers have reviewed the AI4PDEs, including DL for PINN and related parameter identification problems 95, promising directions of AI4PDEs 96, and AI4PDEs in computational mechanics 97. Function learning approximates the mapping between finite-dimensional tensor spaces, while operator learning approximates the mapping between infinite-dimensional function spaces. A representative operator is the solution operator of PDEs, which maps the IC/BCs or parameters (e.g., operating, geometrical, physical) to the solution functions. Function learning can only learn a single PDE solution per training, so the model needs to be retrained once conditions or parameters change. As a progress, operator learning can learn parametric PDE solutions at once, providing a method for rapid PDE-based predictions in practice. When the model of operator learning is an NN, it is referred to as neural operator (NO). Two common NOs are the deep operator network (DeepONet) 98 and the Fourier neural operator (FNO) 99. However, they are both purely data-driven, so their physics-informed versions have also been proposed, i.e., the PI-DeepONet 100 and PI-FNO 101.
@@
Based on these reviews, some conclusions can be drawn. (1) The scope of PIML has been expanded from the classical PINNs to a broader sense. Table 1 compares the PIML in a narrow and in a broad sense from four aspects: model architecture, ways to inform physics, form of physics, and differentiation. In this paper, both types of PIML will be introduced. Unless otherwise specified, PIML in this paper refers to PIML in a broad sense, while PINNs refer to PIML using NNs as the representation model. Besides, some related terms, such as "physics-guided", "physics-embedded", "physics-encoded", "physics-enforced", "physics-constrained", and "physics-preserved", will all be viewed as "physics-informed" in a broad sense in this paper. (2) Many reviews have consensus on three ways to inform physics 68, 83, 85, 89, 91: through data, (ML) model, and optimization, as summarized in Table 2. Although the data is obtained by scientific means, ML based solely on physics-informed data is still purely data-driven, so it is a weak mechanism for informing physics, as stated in Reference 68. Therefore, this paper mainly introduces the latter two methods, and the first way will be introduced only if it has a strong embedding. (3) There is a lack of systematic reviews of PIML for combustion, which is the aim of this paper. (4) Many reviews introduce PIML and NOs together. However, most NOs are purely data-driven. This paper will only introduce PIML and physics-informed NOs (PINOs) in detail. By our definition, the latter can also be regarded as a kind of PIML.
@@
Model architecture MLP All models: CNN 102, 103, Transformer 104, DeepONet 100, 105, 106, 107, 108, 109, 110, 111, KAN 112, 113
How to inform physics Loss function All ways: Model architecture 76, 77, 78, input transformation, output correction 114, 115, 116, 54, 55, numerical solver 117, 118, 116, 119
Form of physics PDEs All forms: algebraic equation, inequality, known coefficients 120, 121, 122, 123
Differentiation AD All methods, such as ND: convolution kernels for space 102, 103, classical schemes for time 124, 76
@@ 2.3 Developments
This subsection will briefly introduce the developments of PIML, focusing mainly on PINNs that use ODE/PDEs as physical constraints, as this is the main and classical part of PIML. Most introductions are brief; for detailed introduction, please refer to the reviews in Section 2.2. Consistent with the taxonomy of References 87, 88, this paper divides PIML into three components: representation model, optimization process, and PDE system, which will be introduced separately in this subsection. Then some miscellaneous important topics will be introduced.
@@ 2.3.1 Representation model
Activation function. Activation functions ($\sigma$ in Figure 4) introduce nonlinearity to an NN. In order to implement AD to solve PDEs, the activation functions of PINNs need to have good differentiability. The $\tanh()$ and $\sin()$ are two common choices. Some adaptive activation functions have been proposed to accelerate the PINNs 137, 138.
@@
Architecture. Based on the representation way of the NN, PINNs can be divided into two categories: continuous and discrete PINNs, which are compared in Table 3. Continuous PINNs follow the coordinate-to-variable (query-to-value) mapping way, i.e., $\widehat{\mathbf{u}}=\mathrm{NN}(\mathbf{z})$, while discrete PINNs follow the field-to-field mapping way, i.e., $\widehat{\mathbf{U}}_{\mathrm{out}}=\mathrm{NN}(\mathbf{U}_{\mathrm{in}})$, where $\mathbf{U}_{\mathrm{in}}$ and $\widehat{\mathbf{U}}_{\mathrm{out}}$ are the input and output fields of the NN, respectively. Continuous PINNs provide continuous representations because they can directly predict the field values at any given query coordinates without interpolation, while discrete PINNs provide discrete representations because $\mathbf{U}_{\mathrm{in}}$ and $\widehat{\mathbf{U}}_{\mathrm{out}}$ are always discrete fields on meshes. When $\mathbf{U}_{\mathrm{in}}$ and $\widehat{\mathbf{U}}_{\mathrm{out}}$ are the fields defined on coarse and fine grids, respectively, the task is called super-resolution (SR) and has been applied to combustion (see Section 5.5). When $\mathbf{U}_{\mathrm{in}}$ and $\widehat{\mathbf{U}}_{\mathrm{out}}$ are fields at time $t$ and $t+\Delta t$, respectively, the task is usually referred to as autoregressive (AR) learning 139, 140. In this case, the NN can be called a time stepper, next-step predictor, forward propagator 141, or emulator 141, 140. On the contrary, the classical continuous PINNs belong to the paradigm of non-autoregressive (NAR) learning 139. The AR paradigm has been widely used in surrogate modeling of chemical kinetics. See Section 4.2 for details. Note that the AR and NAR learning can also be combined to form a hybrid paradigm 139, 142, 143, with DeepONet being a typical method. Instead of using AD in continuous PINNs, discrete PINNs employ numerical differentiation (ND) to compute derivatives. In other words, the PDE residual is $\mathcal{F}[\widehat{\mathbf{U}}_{\mathrm{out}}(\mathbf{U}_{\mathrm{in}})]$ computed on $\mathbf{U}$ using ND, instead of $\mathcal{F}[\widehat{\mathbf{u}}(\mathbf{z})]$ in Equation (2.5) computed on $\widehat{\mathbf{u}}$ and $\mathbf{z}$ using AD. For spatial derivatives, ND can be achieved using differential convolution kernels (similar to the gradient operators in computer vision 144), so convolutional NNs (CNNs) are usually adopted. For temporal derivatives, ND can be achieved via classical numerical schemes, such as the Euler scheme. In this case, recurrent NNs (RNN) can be used for modeling the temporal evolution. The choice between continuous and discrete PINNs will be discussed in Section 8.2.
@@
  • Continuous PINNs. The architectures of continuous PINNs mainly include the MLP, ELM 145, radial basis network (RBN) 146, PointNet 147, DeepONet 98, and Kolmogorov-Arnold network (KAN) 148, 149. Their physics-informed versions have been proposed: PI-ELM 150, PI-RBN 151, PI-PointNet 152, PI-DeepONet 100, and PI-KAN 112, 113. Moreover, the basic MLP has been modified in PINNs, including the multiple MLPs 153, modified MLP 82, and the physics-informed residual adaptive networks (PirateNets) 154, which tackles the issue that PINNs have difficulty using very deep NNs. Some generative models, such as the generative adversarial network (GAN) 155 and variational autoencoder (VAE) 156, can also use MLPs as the building blocks and be extended to the PI-GAN 157 and PI-VAE 158.
@@
  • Discrete PINNs. Basic architectures of discrete PINNs include the RNN (represented by the long short-term memory (LSTM) network 159), CNN 160, 24, recurrent CNN (RCNN), GNN 161, and Transformer 162. Some generative models usually use them as the backbone, such as the VAE, GAN, and diffusion model (DM) 163, 164, 165. The basic architectures can also serve as the building blocks of NOs, such as the Fourier neural operator (FNO) 99. Most of these NNs have their physics-informed extensions, including the PI-LSTM 166, PI-CNN 102, 103, PI-RCNN (or PhyCRNet) 124, PI-GNN 167, PI-Transformer 168, 104, PI-DM 169, and PI-FNO 101. Note that the architectures mentioned in continuous PINNs can also be used in discrete PINNs as long as the mapping is a field-to-field way. The same goes for the generative models, which are architecture-independent. See examples in Section 4.2 and Section 5.5. In fact, the discrete time models proposed in the original paper of PINN can be viewed as a discrete PINN on the temporal dimension 69, which adopted the MLP and Runge-Kutta scheme. However, the architectures just mentioned (such as the CNNs) are rarely applied to continuous PINNs because they typically do not take coordinates as network inputs.